# LongAnimation: Long Animation Generation with Dynamic Global-Local Memory

### v1

@@ 2.1 Long Video Generation
Visual generation has gradually evolved from controllable T2I [19, 13, 50] to controllable T2V [36, 22, 24, 23, 21], in which a key issue is long video generation. Training-free studies [28, 20] extend the video length by window fusion or frequency control. DitCTrl [2] follows window fusion method in DiT [27] architecture. Training-based studies [30, 8, 48] usually rely on the most recent segment to generate the next segment. The most recent study [11] dynamically updates historical segment features by fine-tuning LoRA during inference, achieving longer memory. However, long video generation with finer-grained control (e.g., sketch, ID) is more challenging as it requires more precise long-term consistency of control conditions, not just with text. Some studies [12, 49] introduce local fusion but rarely consider global consistency. LongAnimation innovatively uses Long Video Understanding model to extract global temporal features for long controllable video generation without fine-tuning during inference.
@@ 2.2 Long Video Understanding
The multimodal community has explored many methods for long video understanding (LVU). Early studies [18, 16, 10] processes videos through sparse sampling (e.g., 16 and 24 frames). However, sparse sampling discards numerous frames, causing biased understanding. To achieve better understanding, some studies [38, 31, 37, 17, 41, 6] compress video features to feed more frames into LVU models. Some studies [17, 41, 6] reduce the token numbers per frame or use token merging to compress features. Recent Video-XL [31] and ReTAKE [37] use KV cache to compress visual input, achieving fine-grained visual perception. We use Video-XL, which can process thousands of frames, to extract fine-grained global historical features.
@@ 3 Methodology
The pipeline of LongAnimation is depicted in Fig. 2, which mainly consists of three parts: SketchDiT, Dynamic Global-Local Memory and Color Consistency Reward. During training, given a reference image $I$, $F$ frames of sketches $\{S_{f}\}_{f=1}^{F}$, and corresponding text descriptions, LongAnimation extracts hybrid reference features of these inputs through SketchDiT, which is designed to facilitate the subsequent Dynamic Global-Local Memory mechanism. When generating subsequent segments, Dynamic Global-Local Memory dynamically compresses historical features, which adaptively extracts global consistency features related to the current hybrid features obtained from SketchDiT, ultimately generating $F$ frames of long-consistency animation $\{I_{f}\}_{f=1}^{F}$. Color Consistency Reward is used to refine long-term color consistency during training while color consistency fusion is used to smooth transition segments during inference. In this section, we will first introduce preliminaries. Then, we will describe the SketchDiT, Dynamic Global-Local Memory and Color Consistency Reward. Next, we will introduce the color consistency fusion.
@@ 4.1 Experimental Setups
Implementation. Our model is implemented on CogVideoX-1.5-5B [46], which can stably generate 81 frames. Sakuga-42M [26] is used as our training dataset. Our work is solely for non-commercial scientific research. We filter high-aesthetic and dynamic video clips with lengths greater than 91 frames, retaining about 80k videos for training. The model is trained on 6 A100 GPUs with a learning rate 1e-5. The model is trained at a resolution of $1024\times 576$. We set the SketchDiT layers $L=6$ while CogVideoX-1.5 layers $N=42$. SketchDiT is trained for 30k steps in the first stage, and GLM is trained for 10k steps in the second stage. Finally, we use Color Consistency Reward to refine the coloring ability for 10k steps further. More details are in Appendix A.1.
Test Dataset. We randomly select 3k samples from the Sakuga test set for short-term coloring testing. To ensure a fair comparison with past models, we only use the first 14 frames generated by each model (42k frames in total) for short-term analysis. For long-term animation generation, we select videos with more than 300 frames from the Sakuga test and filter samples with aesthetic scores above 0.8 and dynamic scores above 0.4 (200 videos, approximately 100k frames) to test long-term generation.

### v2

@@ 2.1 Long Video Generation
Visual generation has gradually evolved from controllable T2I [19, 13, 50] to controllable T2V [36, 22, 24, 23, 21], in which a key issue is long video generation. Training-free studies [28, 20] extend the length by window fusion or frequency control. DitCTrl [2] follows window fusion in DiT [27] architecture. Training-based studies [30, 8, 48] usually rely on the most recent segment to generate the next segment. The most recent study [11] dynamically updates historical segment features by fine-tuning LoRA during inference. However, long video generation with finer-grained control (e.g., sketch, ID) is more challenging as it requires more precise long-term consistency of control conditions. Some studies [12, 49] introduce local fusion but rarely consider global consistency. LongAnimation innovatively uses Long Video Understanding model to extract global temporal features for long controllable video generation.
@@ 2.2 Long Video Understanding
The multimodal community has explored many methods for long video understanding (LVU). Early studies [18, 16, 10] processes videos through sparse sampling (e.g., 16 and 24 frames). However, sparse sampling discards numerous frames, causing biased understanding. To achieve better understanding, some studies [38, 31, 37, 17, 41, 6] compress video features to feed more frames into LVU models. Some studies [17, 41, 6] reduce the token numbers per frame or use token merging to compress features. Recent Video-XL [31] and ReTAKE [37] use KV cache to compress visual input, achieving fine-grained visual perception.
@@ 3 Methodology
The pipeline of LongAnimation is depicted in Fig. 2, which mainly consists of three parts: SketchDiT, Dynamic Global-Local Memory and Color Consistency Reward. During training, given a reference image $I$, $F$ frames of sketches $\{S_{f}\}_{f=1}^{F}$, and corresponding text descriptions, LongAnimation extracts hybrid reference features of these inputs through SketchDiT, which is designed to facilitate the subsequent Dynamic Global-Local Memory mechanism. When generating subsequent segments, Dynamic Global-Local Memory dynamically compresses historical features, which adaptively extracts global consistency features related to the current hybrid features obtained from SketchDiT, ultimately generating $F$ frames of long-consistency animation $\{I_{f}\}_{f=1}^{F}$. Color Consistency Reward is used to refine long-term color consistency during training while color consistency fusion is used to smooth transition during inference. In this section, we will first introduce preliminaries. Then, we will describe the SketchDiT, Dynamic Global-Local Memory and Color Consistency Reward. Next, we will introduce the color consistency fusion.
@@ 4.1 Experimental Setups
Implementation. Our model is implemented on CogVideoX-1.5-5B [46], which can stably generate 81 frames. Sakuga-42M [26] is used as our training dataset. Our work is solely for non-commercial scientific research. We filter high-aesthetic and dynamic video clips with lengths greater than 91 frames, retaining about 80k videos for training. The model is trained on 6 A100 GPUs with a learning rate 1e-5. The model is trained at a resolution of $1024\times 576$. We set the SketchDiT layers $L=6$ while CogVideoX-1.5 layers $N=42$. SketchDiT is trained for 30k steps in the first stage, and GLM is trained for 10k steps in the second stage. Finally, we use Color Consistency Reward to refine the coloring ability for 10k steps further.
Test Dataset. We randomly select 3k samples from the Sakuga testset for short-term coloring testing. To ensure a fair comparison with past models, we only use the first 14 frames generated by each model for short-term analysis. For long-term generation, we randomly selected 200 videos, averaging 500 frames to test long-term generation.
@@ 5 Conclusion
6 Acknowledgement
This research is supported by Artifcial IntelligenceNational Science and Technology Major Project 2023ZD0121200, National Natural Science Foundation of China under Grant 62222212 and 623B2094.